OpenStudy (anonymous):

Write each expression as the product of two binomials? 1. 3b(b-2) - 4(b-2) 2. 2x(y+4) + 3(y+4)

Vocaloid (vocaloid):

hint: A(B) - C(B) = (A-C)(B) A(B) + C(B) = (A+C)(B)